# LLMs, You Can Evaluate It! Design of Multi-perspective Report Evaluation for Security Operation Centers

### v1

@@ Abstract
Security operation centers (SOCs) often produce analysis reports on security incidents, and large language models (LLMs) will likely be used for this task in the near future. We postulate that a better understanding of how veteran analysts evaluate reports, including their feedback, can help produce analysis reports in SOCs. In this paper, we aim to leverage LLMs for analysis reports. To this end, we first construct a Analyst-wise checklist to reflect SOC practitioners’ opinions for analysis report evaluation through literature review and user study with SOC practitioners. Next, we design a novel LLM-based conceptual framework, named MESSALA, by further introducing two new techniques, granularization guideline and multi-perspective evaluation. MESSALA can maximize report evaluation and provide feedback on veteran SOC practitioners’ perceptions. When we conduct extensive experiments with MESSALA, the evaluation results by MESSALA are the closest to those of veteran SOC practitioners compared with the existing LLM-based methods. We then show two key insights. We also conduct qualitative analysis with MESSALA, and then identify that MESSALA can provide actionable items that are necessary for improving analysis reports.
@@ 1 Introduction
Cyberattacks have increased across many organizations, and establishing a security operation center (SOC) to analyze security alerts and provide responses has become urgent and crucial for each organization in recent years. An important mission for SOCs is, in addition to providing security alerts and their analyses, to write analysis reports with clear and actionable insights into their underlying cyberattacks [sharma2024decoding]. Interestingly, in proportion to the quality of analysis reports, various stakeholders, including not only analysts but also their managers, can understand the content of security alerts, and hence, the entire organization will be aware of the security [jawad2024m].
However, writing analysis reports for security alerts requires SOC practitioners to force a heavy workload, because these reports must contain actionable insights [kokulu2019MatchedandMismatchedSOCs]. To reduce this workload, methods to extract actionable insights for cybersecurity are in high demand [albanese2025towardsAI-driven]. Although there are several tools to generate analysis reports for security alerts [loumachi2025advancing, zhao2025information], automatically generated reports often contain inaccurate information, thereby affecting the SOC performance [kokulu2019MatchedandMismatchedSOCs]. Even when the state-of-the-art large language models (LLMs) are utilized, they cause hallucinations [CHEN2024SurveyoflargelanguagemodelsforCyberThreat, hasanov2024applicationoflargelanguagemodels]. Moreover, writing analysis reports to the exact requirements of report writing guidelines is stressful [nepal2024burnout]. The reason is that evaluation of analysis reports should identify any lack of descriptions as well as judgments from the perspectives of experts [RYAN2012Quantifying]: for example, “whether response is actionable from the content." The most common mitigation approach for the above problem is to constantly tune and adjust reports [kokulu2019MatchedandMismatchedSOCs], while the evaluation of analysis reports for SOCs needs expertise knowledge [kerstenfield2025tier1analyst].
Based on the above background, in this paper, we discuss the evaluation of analysis reports for security alerts in SOCs with LLMs in order to improve the quality of the reports. Specifically, we discuss the following research questions in this paper:
@@
What are the evaluation criteria that guarantee the quality of analysis reports from expertise knowledge?
@@
Can LLMs quantitatively evaluate the quality of analysis reports in accordance with judgment from the perspective of experts under the defined evaluation criteria?
@@
Can LLMs qualitatively evaluate the quality of analysis reports with feedback for judgment from the perspective of experts under the defined evaluation criteria?
We first propose Analyst-wise checklist as the evaluation criteria to reflect SOC practitioners’ knowledge for the analysis report as the answer to RQ1. To design this, we also conducted a literature review and a user study with semi-structured interviews. We collected 13 public guidelines and handbooks in IT and OT domains as thematic analysis in the literature review, and contacted 15 SOC practitioners with diverse backgrounds in these domains as template analysis [bushra202299False]. We also conduct a validation test of the checklist with multiple SOC practitioners. (See Section 3 for details.)
Next, under on the above checklist, we propose a novel framework, called Multi-perspective Evaluation System for Security Analysis using Llm Assistance (MESSALA), to maximize the evaluation of analysis reports in SOCs. In a nutshell, MESSALA guides LLMs in evaluating analysis reports by leveraging the Analyst-wise checklist from multiple perspectives. As described in Section 4, MESSALA can imitate human cognition to obtain expert knowledge by the checklist from the perspective of human experts, and therefore, can return appropriate evaluation results with feedback to the given analysis reports. When we conduct extensive experiments with MESSALA, we identify that MESSALA outperforms the existing methods [fu2023gptscore, liu2023g-eval] in quantitative evaluation. Thus, as the answer to RQ2, we confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports from the perspective of human experts. (See Section 5 for details.)
Third, we examine whether the feedback returned from MESSALA aligns with judgments from the perspective of human experts through qualitative evaluation. According to the multi-metric rating, MESSALA obtains higher ratings in most metrics by virtue of feedback, which is more specific and easier to understand than common LLM baselines. It means that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback. (See Section 6 for details.)
@@
  • We design an Analyst-wise checklist that contains items for analysis report evaluation in order to reflect SOC practitioners’ knowledge through literature review and semi-structured interviews as the answer to RQ1.
  • We propose MESSALA, a novel framework that maximizes report evaluation and provides feedback from the perspective of SOC practitioners, and confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports through extensive experiments. This is identical to the answer to RQ2.
  • we examine the feedback returned from MESSALA, and then identify that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback, which is the answer to RQ3.
@@ 2 Related Work and Our Problem Setting
2.0.1 Large-Language Models
LLMs are models that combine language processing with data generation. They learn observed tokens as sequences of characters and then find the most probable tokens in the output sequences [sandoval2023LostAtC]. Input to an LLM is called a prompt. LLMs can provide comparable performance to a veteran data analyst [cheng2023-IsGPTaGoodDataAnalyst], and are used to evaluate data as LLM-as-a-judge services [hu2024llm]. We focus on GPTScore [fu2023gptscore] and G-Eval [liu2023g-eval]. A common problem for LLMs is hallucinations with incorrect outputs [huang2025hallucination].
LLMs have been utilized in many cybersecurity research, such as data collection [BAYER2023multi-levelfine-tuning, papa2023AnAttackersDream, sharma2023HowWellDoesGPTPhishPeople], data pre-processing [liu2023NottheEndofStory, wadhwa2023RevisitingRelationExtraction], feature extraction [BOFFA2024Logprecis, jiang2024Xpert],attack detection [ali2023HuntGPT, Ghourabi2023EnhancingSpamMessage, koide2024ChatPhishDetector]and report generation [Gupta2023fromChatGPTtoThreatGPT, mitra2024LocalIntel, perrina2023agir, yamin2024applications]. However, LLMs require further fine-tuning for each specific task of cybersecurity [ferrag2023RecolutionizingCyberThreat]. We discuss a report evaluation method based on LLMs for cybersecurity.
@@ 2.0.2 Security Operation Centers
SOCs address various aspects of cyberattacks, including data collection, determining the presence of intrusions, and responding to threats [Tilbury2024HumansandAutomation]. The major role of SOCs is to detect cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible], investigate security alerts [yen2013Beehive, zhong2019LearningfromExpertsExperimence], and make judgments on these alerts [kokulu2019MatchedandMismatchedSOCs]. SOCs roughly consist of two processes, i.e., the evaluation process by analysts and the judgment process by managers. To support and connect both processes, creation of analysis reports is important [kokulu2019MatchedandMismatchedSOCs].
Although SOCs in recent years have introduced automation tools [AGYEPONG2023ASystematicMethod, shahjee2022IntegratedNetwork] to reduce workload [Tilbury2024HumansandAutomation], it is still heavy [nepal2024burnout, kerstenfield2025tier1analyst, kersten2024securityalert]. Analysts in SOCs also need to understand cyberattacks based on fragmented information [VANDERKLEIJ2022DevelopingDecisionSupport] and frequently require source information to determine appropriate responses [hassan2019NoDose, happa2021Assessing]. Although machine learning, including LLMs, has been developed for analysts in SOCs in recent years [ANDRADE2019CognitiveSecurity, ban2021CombatSecurityAlert, baroni2021Self-AwareEffective, gonzalez2021SIEM, HUSAK2022Crusoe, mitra2024LocalIntel, jiang2024Xpert, albanese2025towardsAI-driven, singh2025llms, cui2025trafficllm, li2024dollm, singh2025contextbuddy], there are limitations for providing analysis reports [michelet2024chatgpt] and security advice [chen2023can] from LLMs.
@@ 2.0.3 Evaluation of Analysis Reports in SOCs
We focus on evaluation of analysis reports in SOCs as the main problem setting. The current report generation tools have limitations since LLMs are often unable to generate analysis reports with high quality [michelet2024chatgpt]. Instead, we aim to explore whether LLMs can effectively evaluate these reports. We note that the evaluation of analysis report remains challenging: general LLMs may fail to align with human judgment because these reports often include domain-specific contexts [chiang2023closer].
@@
We aim to design a system that takes analysis reports as input and returns evaluation results as output. The output includes both quantitative scores and qualitative feedback from the perspectives of SOC analysts, such as actionable feedback. We consider that the readers of analysis reports are both analysts and managers, while the generation of the reports themselves are outside the scope of this paper.
@@ 3.2 Collection of Candidate Items through Literature Review
Below, we present the set of items that should be documented in an analysis report and the categories to which they belong.

### v2

# LLMs, You Can Evaluate It! Design of Multi-perspective Evaluation for Reports in Security Operation Centers

@@ Abstract
Security operation centers (SOCs) often produce analysis reports on security alerts, and large language models (LLMs) will likely be used for this task. While producing analysis reports with actionable insights requires constantly evaluating the reports of SOC practitioners, it is unclear whether LLMs appropriately evaluate the reports due to a lack of evaluation criteria. In this paper, we discuss evaluations using LLMs for analysis reports. To this end, we first design Analyst-wise Checklist that contains SOC practitioners’ evaluation criteria for analysis reports in the real world through literature review and interviews with SOC practitioners as a user study. Next, we design a novel LLM-based multi-perspective evaluation framework, named MESSALA, that maximizes report evaluation and provides feedback with analysis reports on veteran SOC practitioners’ perceptions by leveraging the checklist above. We also conduct quantitative evaluation and qualitative evaluation with real-world analysis reports. For the quantitative evaluation, we demonstrate that MESSALA can appropriately rate each analysis report on the evaluation criteria from the perspective of veteran SOC practitioners compared with existing LLM-based methods. For the qualitative evaluation, we identify that MESSALA can provide feedback comments that are actionable for improving analysis reports under the evaluation criteria.
@@ 1 Introduction
Cyberattacks have increased across many organizations, and establishing a security operation center (SOC) to analyze security alerts and provide their responses has become urgent and crucial for each organization in recent years. An important mission for SOCs is, in addition to providing security alerts and their responses, to write analysis reports with clear and actionable insights into their underlying cyberattacks [sharma2024decoding]. Interestingly, in proportion to the quality of analysis reports, various SOC practitioners, including not only analysts but also their managers, can understand the content of security alerts, and hence, the entire organization will be aware of cyberattacks [jawad2024m].
However, writing analysis reports for security alerts forces a heavy workload on SOC practitioners, because these reports must contain actionable insights [kokulu2019MatchedandMismatchedSOCs]. To reduce this workload, tools, such as large language models (LLMs), to introduce actionable insights into analysis reports are in high demand [albanese2025towardsAI-driven]. Although there are several tools to generate analysis reports for security alerts [loumachi2025advancing, michelet2024chatgpt], automatically generated reports often contain inaccurate information, which may degrade performance in SOCs [kokulu2019MatchedandMismatchedSOCs]. Even when the state-of-the-art LLMs are utilized, they often cause hallucinations such as inaccurate information [huang2025hallucination]. As another motivation, writing analysis reports to the exact requirements of report-writing guidelines is stressful [nepal2024burnout]. The reason is that evaluation of analysis reports should identify any lack of descriptions as well as judgments from the perspectives of veteran SOC practitioners [RYAN2012Quantifying]. Although the most common mitigation approach for the above stress is to constantly evaluate analysis reports [kokulu2019MatchedandMismatchedSOCs], evaluations of the reports need evaluation criteria based on expert knowledge [kerstenfield2025tier1analyst, kersten2023give]. Moreover, the current guidelines are often abstract and may fail to capture the evaluation criteria from the perspective of veteran SOC practitioners [johnson2016guide, agyepong2020towards]. Namely, even evaluations of analysis reports using LLMs are challenging due to a lack of enough evaluation criteria.
Based on the above background, in this paper, we discuss evaluations for analysis reports in SOCs using LLMs in order to improve the quality of the reports. Specifically, we discuss the following research questions in this paper:
@@
What are the evaluation criteria that guarantee the quality of analysis reports from SOC practitioners’ knowledge?
@@
Can LLMs quantitatively evaluate the quality of analysis reports from the perspective of SOC practitioners under the defined evaluation criteria?
@@
Can LLMs qualitatively evaluate the quality of analysis reports with feedback from the perspective of SOC practitioners under the defined evaluation criteria?
We first design the Analyst-wise Checklist, which contains the evaluation criteria in its specific items to reflect SOC practitioners’ knowledge for evaluating analysis reports in the real world, as the answer to RQ1. To design this, we also conducted a literature review of 13 public guidelines and handbooks in the IT and OT domains and then semi-structured interviews with 15 SOC practitioners. We also conducted a validation test to confirm the checklist with multiple SOC practitioners. (See Section 3 for details.)
Next, based on the above checklist, we propose a novel framework, called Multi-perspective Evaluation System for Security Analysis using Llm Assistance (MESSALA), to maximize the evaluation of analysis reports in SOCs. In a nutshell, MESSALA guides LLMs in evaluating analysis reports by leveraging the Analyst-wise Checklist from multiple perspectives. As described in Section 4, MESSALA can imitate SOC practitioners’ cognition [zhong2016automate] to obtain expert knowledge from the checklist, and therefore, can return appropriate evaluation results with actionable feedback for each analysis report. When we conduct quantitative evaluations with MESSALA, we identify that it can rate analysis reports on the evaluation criteria compared with the existing methods [fu2023gptscore, liu2023g-eval]. Thus, to answer RQ2, we confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports from the perspective of veteran SOC practitioners. (See Section 5 for details.)
Third, we conduct qualitative evaluations to identify whether the feedback generated by MESSALA aligns with judgments from the perspective of veteran SOC practitioners. As a result, we confirm if MESSALA can provide more actionable feedback with both novice and veteran SOC practitioners than the existing methods, because it is more specific and easier to understand. It means that MESSALA also enables LLMs to qualitatively evaluate analysis reports through its feedback. (See Section 6 for details.)
@@
  • We design the Analyst-wise Checklist that contains evaluation criteria for evaluating analysis reports to reflect real-world SOC practitioners’ knowledge through literature review and interviews with SOC practitioners as the answer to RQ1.
  • We propose MESSALA, a novel framework that maximizes evaluation for each criterion and provides feedback from the perspective of SOC practitioners, and confirm if MESSALA enables LLMs to quantitatively evaluate analysis reports under the evaluation criteria. This is identical to the answer to RQ2.
  • We qualitatively evaluate the feedback returned from MESSALA, and then identify that MESSALA enables LLMs to qualitatively evaluate analysis reports through its feedback under the evaluation criteria, which is the answer to RQ3.
@@ 2 Related Work and Our Problem Setting
2.0.1 Large Language Models
LLMs generate text outputs from natural-language prompts through probability distributions of tokens. Various use cases have been proposed; for example, LLMs provide performance comparable to that of veteran data analysts [cheng2023-IsGPTaGoodDataAnalyst] and can be used as LLM-as-a-judge services for data evaluation [kim2023prometheus]. We focus on GPTScore [fu2023gptscore] and G-Eval [liu2023g-eval]. A common problem for LLMs is hallucinations [huang2025hallucination].
LLMs have been utilized in many cybersecurity research, such as data collection [BAYER2023multi-levelfine-tuning, shibli2024AbuseGPT], attack and threat analysis tasks [gadyatskaya2023ChatGPTKnowsYourAttacks, BOFFA2024Logprecis, shibli2024AbuseGPT], and report generation tasks [Gupta2023fromChatGPTtoThreatGPT, mitra2024LocalIntel, perrina2023agir, loumachi2025advancing, wickramasekara2024framework]. However, the above works do not provide a detailed discussion for evaluations of analysis reports. We focus on the evaluation of reports in this paper.
@@ 2.0.2 Security Operation Centers
SOCs address various aspects of cyberattacks, including data collection, determining the presence of intrusions, and responding to threats [Tilbury2024HumansandAutomation]. The current SOCs have monitors industrial control systems, such as factories, as well as the conventional IT systems [albasheer2022Cyber-attack, oba2024ScoreAndYouShallFind]. The major role of SOCs is to detect cyberattacks [gupta2019AutomatedEventPrioritization, renners2019AdaptiveandIntelligible], investigate security alerts [kerstenfield2025tier1analyst, yen2013Beehive] , and make judgments on these alerts [kokulu2019MatchedandMismatchedSOCs]. SOCs roughly consist of two processes, i.e., the evaluation process by analysts and the judgment process by managers. To support and connect both processes, analysis reports are significantly important [kokulu2019MatchedandMismatchedSOCs].
Although SOCs in recent years have introduced automation tools [singh2025contextbuddy, turcotte2025automated, shahjee2022IntegratedNetwork, AGYEPONG2023ASystematicMethod] to reduce workload [Tilbury2024HumansandAutomation], it is still heavy [nepal2024burnout, kerstenfield2025tier1analyst, kersten2024securityalert].Analysts in SOCs also need to understand cyberattacks based on fragmented information [VANDERKLEIJ2022DevelopingDecisionSupport] and frequently require source information to determine appropriate responses [happa2021Assessing]. Although machine learning, such as LLMs, has been developed for analysis in SOCs recently [ANDRADE2019CognitiveSecurity, gonzalez2021SIEM, HUSAK2022Crusoe, mitra2024LocalIntel, albanese2025towardsAI-driven, singh2025contextbuddy], there are limitations for providing analysis reports [michelet2024chatgpt, loumachi2025advancing, wickramasekara2024framework] and security advice [chen2023can] from LLMs.
@@ 2.0.3 Evaluation of Analysis Reports in SOCs
We focus on evaluations of analysis reports in SOCs as the main problem setting. The current report generation tools [loumachi2025advancing, michelet2024chatgpt, wickramasekara2024framework] have limitations due to hallucinations, since LLMs often cause hallucinations and these tools also lack explicit mechanisms for evaluating of analysis reports. Instead, we aim to explore whether LLMs can appropriately evaluate these reports. We note that the evaluation of analysis reports remains challenging: general LLMs may fail to align with human judgment because these reports include domain-specific contexts [chiang2023closer].
@@
We aim to design an evaluation framework that takes analysis reports as input and returns evaluation results as output. The output includes both quantitative scores and qualitative feedback from the perspectives of SOC practitioners, such as veteran analysts. We consider that readers of analysis reports are both analysts and managers, while the generation of the reports is outside the scope of this paper.
@@ 3.2 Collection of Candidate Items through Literature Review
We present the set of items that should be documented in an analysis report their categories below.